You are here: Home / Documents / Minutes of CASU Meetings / minutes_090817.txt

Plain Text icon — Plain Text, 13 KB (14052 bytes)

File contents

Monday 17th August 2009   APM Meeting tables    11:30-12:30

Present:        MJI, RWA, NAW, EGS, MR
Apologies:      DWE, RGM, JRL, SC, STH

1.  Actions from last meeting
2.  Comments on WFAU minutes
3.  Recent and upcoming meetings
4.  Data archives update     - AAT, ING, WFCAM, UKIRT
5.  Optical/NIR processing   - HAWKI, INT WFC, MegaCam, Subaru, VST .....
6.  WFCAM update
7.  VISTA update
8.  AOB - grant status etc..


1. Actions from the last meeting

MJI  stage I software release on hold due to other pressures - ongoing     <<<<

MJI  ditto technical paper - ongoing                                       <<<<

MJI  ditto utility to convert binary tables plus corrections - ongoing     <<<<

STH  guide to whereabouts of on-line data particularly for ING archive     <<<<
     - ongoing we think                                                

MJI  Petrosian Fix has been done and incorporated in all the catalogues
     held here.  The code has been sent to WFAU to do the same with the
     catalogues at their end, plus form basis for direct DB update

MJI  generated an overview summary plot of sky brightness levels of recent 
     VISTA on-sky data, plus the same for photometric zero-points for good
     measure. Both are now available in the revamped VISTA technical pages.
     The average sky background in Ks is close to 13.0, somewhat brighter 
     than for WFCAM in K, but in line with what the ETC and ESO expected
     He also further compared ETC fact with fiction using the latest
     overhead estimates from Thomas Szeifert with satisfactory agreement

RWA  have created an off-line LTO4 backup system for VISTA raw data. The
JRL  VISTA technical pages contain an internal record of which backup LTO4 
     tapes contain which raw VISTA data.

STH  revisit the linearity, zero-points and colour equations for VISTA     <<<<
     - ongoing we hope                                                 

SC   fixed the flatlining apm29 temprature monitoring web page   

EGS  has produced an analysis of the deep stack properties of the VISTA   
     J-band pawprint test data and made the report available through the 
     VISTA technical pages

2. Comment on WFAU minutes

No signfificant comments.  Assorted communications with WFAU and JAC 
highlighted a few issues to do with either screwed-up 09A files (WCS and
missing fluxes - the latter due to a disk problem) which have been fixed 
en passant, or a few examples of much earlier data suffering from badly 
computed offsets in the dither stacking process.  These latter were fixed
by re-running through the latest pipeline offset/stacking part.  The
numbers of these in earlier data is so low, it is easier to fix when
they surface.

3. Recent and upcoming meetings

There was a VDMT meeting on July 15th shortly after the previous CASU
meeting, which dealt mainly with discussions about grant funding.

The VISTA IOT on July 29th was missed but the minutes are in the usual 
place, if anyone is interested.

A reminder that there will be an EGAPS meeting in Manchester on Friday 
September 11th.  NAW plans to drive!%^7* and has offered seats in his car 
for anyone else wishing to go.

Another reminder about the UKIRT workshop to be held on September 14th - 16th 
to celebrate 30 years of astronomy with UKIRT with Richard Ellis giving the 
opening talk. MJI will speak on Local Group galaxy properties in the NIR and
STH on the WTS.

In addition to all the other ongoing reviews, STFC have asked for input
for the "STFC Prioritisation of your project/area".  The proforma has been 
split into five sections representing the main priority areas of STFC: 
Excellence, Economic Impact, Societal Impact, Leadership and Synergies, 
with Excellence as the primary criterion and so weighted more heavily.
So research impact irrelevant then !

4. Data archives

All is proceeding smoothly with the AAO, ING and WFCAM archives, apart from 
some restructuring work on the UKIRT Cass archive which is taking longer 
than anticipated.  However, work is progressing toward getting a final 
version of the UKIRT cass archive sorted. The first step has involved backing
up the original sdf data onto new media as the old DLT tapes are beginning to 
fail.  This has involved transferring quite a lot of files from JAC.
When complet the FITS conversions will then be redone as experience with the 
archive has led to improved methods to make the data more internally 
consistent. This should all be finished by the end of 2009.

NAW asked what will happen to the ING archive. MJI confirmed that it will 
continue to be hosted here providing the additional cost is within the
usual operational noise. This includes a small amount of effort once a month
to ingest and check the transfers, tape backups and occasional leverage of
extra disk storage space.

5. Optical/NIR processing : HAWKI, INT WFC, MegaCam, Subaru, VST .....

EGS noted he was doing a lot of processing for research projects using INT 
and CFH data.  He has noticed that the coefficients of the ZPN polynomial 
distortion on the INT WFC have changed slightly since the commissioning
measurements in all bands, but especially in U, so that the WCS rms errors
go up to ~150 mas from ~100mas in griz and to ~400mas in U. This is not a 
serious problem since when doing the band-merging using fitiso_merge an 
internal astrometric refinement onto the reference band is made when
cross-matching, but it is still worth noting.  Refining the coefficients 
improves the general astrometric precision to ~110mas in griz and ~200mas in U. 
MJI commented that the change may be related to replacing the final element in 
the corrector barrel after an accident cracked the original around 5 years ago.

The INT U filter is a liquid CuS04 variety which may be deteriorating and
causing more optical radial distortion which in turn has exacerbated the 
WCS problem in U.

MJI had an e-mail from Janet Drew about the problems with the VST H-alpha 
filter.  One of the 4 segments (D) is somewhat compromised over 25% of its 
area due to a bandpass shift.  However, as there will be no VST data for the
near future due to the rusty mirror cell incident, a discussion of what
to do about the H-alpha filter has been deferred until the upcoming EGAPS
meeting in September.

Janet will visit sometime towards the end of September with a postdoctoral 
student to discuss IPHAS overall calibration issues.  Paul Groot also wants
to come along to discuss UVEX processing and other related issues.

6. WFCAM update.

Semester 09A processing is complete and remaining July processing is also 
complete apart from one stuck night awaiting JRL's return. The recalcitrant 
night has about 50% missing #4 data, on, off,on, off ..... throughout the 
night, confusing the pipeline at various points.

Semester 09B started on July 27th (sic!) and MR reported that data flow 
and reduction is continuing without problems except for occasional glitches, 
especially at the weekends, when the conversion routine dies at the JAC end.  
It is not clear if these are just some recent flukes or symptomatic of 
something else.  MJI said he would contact AJA and see if there was anything
obvious causing this.                                                      <<<<

A few MSBs have had incorrect RECIPE keywords but advance warning from JAC
has enabled these to be corrected before processing begins.  

Reprocessing proceeds apace with the Petrosian fix finally applied to all 
the earlier catalogues at this end.  Next in the queue are the DR6 GPS 
nebulosity enhanced (re)processing request to be followed by UDS J and 
H-band data from 05B-07B.  NAW said that he had met Phil Lucas at the IAU 
and that Phil was happy with the nebulosity (re)processing. 

MJI said he had had a few recent complaints going back to 07B data about 
occasional poor stacking results.  These manifest themsevles as either double 
or distorted images on one detector, whereas the others are fine.  These
are investigated and dealt with as they arise and so far have always been
caused by incorrect offsets being computed by the pipeline during dither
stacking.  The latest version of the pipeline now computes overall tangent 
point offsets, rather than on a detector-by-detector basis, and this has cured
the problem.

Alastair Edge visited at the end of July to enquire about testing the pawprint
sky method on DXS data, the default is to use the tile method.  It is not
obvious if this will improve DXS results so MJI asked for a test night
to be selected to trial.

MJI commented that he had recently been made aware of a NIR wide field imager
on KPNO 4m (NEWFIRM).  In spite of the rubbish acronym this consists of 4 
Rockwell 2k x 2k Hawaii-II buttable detectors (i.e. similar to WIRCAM on CFH)
and has a fov of around 28 x 28 arcmin.  He is curious to know what, if any,
science is being done using it.

7. VISTA update

Prior to the VISTA IOT, Thomas Szeifert produced a report summarising the 
experiences of a recent 14 night run using VISTA. In most aspects the overall
system appeared to perform well.  There were at the time a few problems
still to be resolved but these were mostly of a minor nature.  It still looks
like SV will take place during September/October though there are concerns
still to be resolved regarding the turnaround time for CASU to receive and 
process the SV data which may necessitate special data transfer measures.

Meanwhile RWA reported that raw VISTA data has been received, verified and
ingested up to July 29th, though processing awaits JRL's return.

The VISTA technical pages have been significantly enhanced and should contain 
a summary of everything we have learned so far. The VISTA user manual is being
written by ESO based on documentation supplied by VDFS and the VPO.  There
are still some gaps in technical detail required for this hence the importance
of keeping the technical web pages here up to date.

STH has produce a provisional report on the SIQ and within notes that in good 
seeing the 50% enclosed energy criterion is being met for most bands, but not 
the 80% figure. MJI thinks this may be related to the assumed underlying
PSF, probably a Moffat profile with beta=2.5, whereas the real profile is 
likely to be much closer to exponential, if other sites and telescope systems
are anything to go by. If the PSF is treated as exponential the 80% figure 
is probably achieved, though further analysis is required to check this.

EGS has been analysing the deep stacking properties for another of the VISTA
technical specs.  He has produced an assortment of deep stacks from a series
of pawprint images on a chosen field obtained during the period May 22-27. 
The analysis is now available on the technical pages and shows that the sky 
noise decreases with squareroot number of images in the stack as expected. 

The ETC for Vista has been updated by MJI using the latest overheads and gives 
plausible numbers for template execution times.

One of the many things PSB did was to set up a VIRCAM helpdesk account here
in Cambridge.  All user queries and ESO tickets should come in via this 
account.  For now only JRL and MJI are automatically forwarded these but
if anyone else is desperate to see them let us know.

The current wish-list of jobs for upgrading the VISTA cataloguing software
stands at: average confidence in aper3 encoded in #55; half-light radius 
analysis to replace Hall measures; elliptical apertures for all galaxian
measures; PSF analysis; Sersic fits.  All we need are people to do it and 
await the grant resolution with bated breath.

8. AOB

Ten new CPU/disk storage units have been delivered for the GAIA project. 
Among other things these will produce on average another 5 kw of heat which  
necessitates an upgrade of the AC system.  Two extra AC units each capable 
of handling 10 kW will be installed, one of which will be set-up, at least 
initially, to act as a backup system.  Another computer rack is also 
required placing more strain on the space available. To make room, an 
alternative storage location is being sought for some of the CD/DVD cabinets.
Quotations for the AC upgrade are in hand as is purchasing of another rack 
to house the new GAIA kit and switched power supplies to deliver the heat.

STH has also acquired a similar CPU/disk storage box for his WTS work which 
has been setup in the other partitioned off computer room.

By cunningly replacing some 400GB SATA disks attached to 3-ware 9500 series
controllers, with 1TB SATA disks MJI, and a visiting JMI, have converted 4TB 
to 10TB at low cost, with no extra space requirements and no extra heat output.
This has been trialled successfully for one of the two disk controller 
partitions in a Chenbro chassis and will be rolled out across others as 
extra bulk storage space is needed.

Some of the spare SATA disks thus liberated have been incorporated in one
of the WFCAM processor units, which after the requisite 3-ware upgrade has 
speeded up i/o on this to a respectable 75 MB/s, some 2-3 times faster than 
before which was based on pre-SATA disks.

Finally, MJI relayed the latest installment in the ongoing grant saga whereby
now 1 rolling grant = 1 updated rolling grant + 1 standard research grant + 
1 other standard grant = what we were asking for over 2 years ago =
what the review panel recommended = ?????? ....... to be continued.

Continuing actions

EGS  add catalogue software and binary table conversion software to software 
MJI  release pages  

MJI  finish off technical paper updates

MJI  write utility to convert binary tables with all corrections

STH  update data storage web page particularly with respect to ING archive

STH  revisit average zero-points and colour equations for VISTA and check
     non-linearity correction is working by cf. with 2MASS data

New actions

MJI  contact AJA and see if there is any obvious issue with the occasional
     faux pas during disk transfers